cube of osmium metal has 1.500 cm on a side and has a mass of 0.07631 Kg

at 25°C. Calculate the density of the cube. 


1
Expert's answer
2021-08-16T02:45:49-0400

Solution:

Volume of cube with side lenghts 's':

V = s × s × s = s3

s = 1.500 cm

Therefore,

V = (1.500 cm)3 = 3.375 cm3


Density = Mass / Volume

Mass = (0.07631 kg) × (1000 g / 1 kg) = 76.31 g

Therefore, the density is:

Density = (76.31 g) / (3.375 cm3) = 22.61 g/cm3

Density = 22.61 g/cm3


Answer: The density of the cube of osmium metal is 22.61 g/cm3